# Some classes satisfying the 2-dimensional Jacobian conjecture

### v3

@@ Abstract.
We construct a non-proper set of two variables polynomial maps and study the nowhere vanishing Jacobian condition of the Jacobian conjecture for this set. We obtain some classes of polynomial maps satisfying the 2-dimensional Jacobian conjecture for both real and complex cases. We prove also that the complex conjecture is true until degree 104.
@@
Jacobian conjecture, polynomial maps, non-proper maps
@@ 1. Introduction
([20, Lemma 2.2]). Then ${\rm det}(J(p,q)(x,y))>0$ for every $(x,y)\in\mathbb{R}^{2}$ since if $t=0$ then $f=y=\frac{1}{x}\neq 0$. However, $(p,q)$ is not automorphism. Our first observation is: by changing variables $t:=t(x,y)$, $h:=h(x,y)$ and $f:=f(x,y)$, we can write the Pinchuk maps $(p,q)$ as polynomial maps of new variables $t,h$ and $f$. The sequence $\{z_{k}\}=\{(k,1/k)\}\subset\mathbb{R}^{2}$ tends to infinity but $t(z_{k})=h(z_{k})=0$ and $f(z_{k})=1/k$. Then the Pinchuk maps $(p,q)$, being polynomial maps of new variables $t,h$ and $f$, satisfy that $(p(z_{k}),q(z_{k}))$ tends to a finite point and therefore they are non-proper and hence, is not automorphism. We call the set of such new variables $\{t,h,f\}$ a set of non-proper variables of the map $F=(p,q)$. By studying the Jacobian condition for polynomial maps of these new variables and classifying them, we obtain some classes that satisfy the Jacobian conjecture for both real and complex cases.
@@
The first part, presented in Section 2, concerns the construction of non-proper variables for both real and complex cases and a classification of non-proper maps written under these non-proper variables. We obtain some explicit classes satisfying the 2-dimensional Jacobian conjecture for both real and complex cases (Corollaries 2.19, 2.21 and 2.22). These results are classified in Remark 2.24. We also provide a possible counter-example for the conjecture (if there exists) in Corollary 2.25.
The second part, presented in Section 3, concerns a particular result to the complex conjecture for degree until 104 (Theorem 3.6). We prove that the complex conjecture is true for degree until 104 using known results studied by Abhyankar [2], Nagata [17, 18], Appelgate and Onishi [3], Magnus [13], Nakai and Baba [16] and Żoła̧dek [24]. It is known that the complex conjecture is true for degree until 100 (Moh [14], 1983). There are some works increasing the degree 100 of Moh, for instance [10], but we emphasize that our proof is a very simple one. There may be a connection between the two parts of the paper for higher degree cases that we intend to study in our next step (Remark 3.8).
@@ 2.1. Construction and examples
We consider $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$ and $F=(F_{1},F_{2}):\mathbb{K}^{2}\to\mathbb{K}^{2}$ a polynomial map of two variables $x$ and $y$. With $z\in\mathbb{K}^{2}$, we denote by $\|z\|$ its Euclidean norm. Let us recall firstly that the polynomial map $F$ is non proper if there is a sequence $\{z_{k}=(x_{k},y_{k})\}_{k\in\mathbb{N}}\subset\mathbb{K}^{2}$ such that $\lim_{k\to\infty}\|z_{k}\|=+\infty$ and $\lim_{k\to\infty}F(z_{k})=(c_{1},c_{2})$, for some finite $c_{1},c_{2}\in\mathbb{K}$. The following remark is important for the definition of non-proper variables sets that we will introduce in Definition 2.2.
@@
In fact, if $F=(F_{1},F_{2})$ is non proper, we can take $u_{0}:=F_{1}$ and $u_{1}:=F_{2}$ and the inverse is obvious. We define:
@@
then $\{u_{0},u_{1},u_{2},u_{3}\}$ is a set of non-proper variables of $F$ since any sequence $\{(1/k,-k+c)\}$, with $c\in\mathbb{K}$, satisfies the first condition in Definition 2.2. The second condition is verified by taking
@@
Now we show that $\{u_{0},u_{1},u_{2},u_{3}\}$ is independent. In fact, for any $z=(x,y)$, we have
@@
In the following example we will show that the functions $t,h,f$ defined in [20] and also in [9] form a set of non-proper variables for the Pinchuk maps and it is independent. Let us remark that we will consider both of real and complex cases, i.e. we consider the polynomials $t,h,f:\mathbb{K}^{2}\to\mathbb{K}$, with $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. In the complex case, it shows that the inverse of Corollary 2.8 is not true.
@@
    In fact, from [20, Lemma 2.1], it is enough to show that there is no $(x,y)\in\mathbb{K}^{2}$ such that
@@
    In fact, as in [20, page 3], set $p=f+h$, then by the Lemma 2.2 in [20], there exists a polynomial $u(f,h)$ such that
@@
As the remark at the end of the paper [20], the Pinchuk map constructed in the proof of [20, Lemma 2.2] has degree 40, where ${\rm deg}(p)=10$ and ${\rm deg}(q)=40$, but one can reduce ${\rm deg}(q)$ to 35 by a suitable choice of the auxiliary polynomial $u(f,h)$. In fact, the polynomial $q$ varies for different Pinchuk map, but always has the form $q=-t^{2}-6th(h+1)-u(f,h)$. In [8] (page 241), van den Essen choose
@@
In [9], basing on Pinchuk’s examples [20], the author provides new counterexamples to the real conjecture but with the possibility to reduce the degree of $q$ to 15: In [9, Corollary 2.6], they prove that if taking $p=f+h$ and
@@
(see the proof of [9, Proposition 2.3]). It follows:
@@ 2.2. A set of non-proper variables
has Newton polygon containing two points (1, 0) and (0, 1) (see Definition of Newton polygon in [2], [17, 18], also in [8, page 93]). In the following we will prove first that for any triple of non-zero natural numbers $(n,r,s)$ and for any polynomial mapping $F:\mathbb{K}^{2}\to\mathbb{K}^{2}$ such that
@@
where $H=(H_{1},H_{2}):\mathbb{K}^{n+1}\to\mathbb{K}^{2}$, then $\{u_{0},u_{1},\dots,u_{n}\}$ is a set of non-proper variables of $F$ (Proposition 2.12 for complex case and Proposition 2.13 for real case). Then we investigate the independence (Definition 2.4) for this set: we consider all the possibilities of triples $(r,s,n)$, we divide them into three cases: $r=1$ (section 2.2.2); $(n=1,\,r\geq 2)$ (section 2.2.3); and $(n\geq 2,r\geq 2)$ (section 2.2.4). We obtain some classes satisfying the Jacobian conjecture for polynomial mappings $F:\mathbb{K}^{2}\to\mathbb{K}^{2}$, $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$ (Corollary 2.21 and Corollary 2.22).
@@
Now if we replace $\mathbb{C}$ by $\mathbb{R}$ in Proposition 2.12, then the sequence $\{(x_{k},y_{k})\}$ chosen in the above proof can be used if $b>0$. Therefore we have:
@@
If $b>0$, then the set of variables $u_{0},u_{1},\ldots,u_{n}:\mathbb{R}^{2}\to\mathbb{R}$ in (2.11) is a set of non-proper variables of any polynomial map $F(x,y)=H(u_{0},u_{,}\ldots,u_{n}):\mathbb{R}^{2}\to\mathbb{R}^{2}$ where $H=(H_{1},H_{2}):\mathbb{R}^{n+1}\to\mathbb{R}^{2}$ is a polynomial map of variables $u_{0},u_{1},\dots,u_{n}$.
@@
Consider the variables $u_{0},u_{1},\dots,u_{n}$ in (2.11).
@@
If $F:\mathbb{R}^{2}\to\mathbb{R}^{2}$ can be written as a polynomial map of variables $u_{0},u_{1},\dots,u_{n}$, i.e. $F(x,y)=H(u_{0},u_{1},\ldots,u_{n})$, where $H=(H_{1},H_{2}):\mathbb{R}^{n+1}\to\mathbb{R}^{2}$ is a polynomial map of variables $u_{0},u_{1},\dots,u_{n}$ in (2.11) with $b>0$, then $F$ is non-proper.
@@
Then any point $z=(x_{0},\sqrt[s]{1/b})$, where $x_{0}\in\mathbb{K}$, is a solution of the system $\dfrac{\partial u_{i}}{\partial x}(z)=0$, for $i=0,1,\ldots,n$ (notice that we consider $b\neq 0$ if $\mathbb{K}=\mathbb{C}$ and $b>0$ if $\mathbb{K}=\mathbb{R}$). ∎
@@
    In the real case $u_{0},u_{1}:\mathbb{R}^{2}\to\mathbb{R}$, the set of non-proper variables $\{u_{0},u_{1}\}$ is dependent with any $r\geq 2$, $s\geq 1$ and $b>0$.
@@
We have ${\partial u_{1}}/{\partial x}=1-rbx^{r-1}y^{s}$. With the same argument as in the proof of Theorem 2.18, the equation ${\partial u_{1}}/{\partial x}=0$ always has a solution in the complex case. In the real case, when $r$ is odd, $s$ is even and $b>0$, the equation ${\partial u_{1}}/{\partial x}=0$ admits al least a solution. ∎
From Proposition 2.6, Corollary 2.19 and Theorem 2.20, we have two explicit classes in which the two-dimensional Jacobian conjecture is true, for complex and real cases, respectively:
@@
$u_{0}=y,\quad u_{1}=x-bx^{r}y^{s}\quad(b\in\mathbb{C}\setminus\{0\},\;r,s\in\mathbb{N}\setminus\{0\}),$
@@
$u_{0}=y,\quad u_{1}=x-bx^{r}y^{s}\quad(b\in\mathbb{R}\setminus\{0\},\;r,s\in\mathbb{N}\setminus\{0\}),$
@@
For a fixed natural number $n\geq 2$, the set of variables
$u_{0},u_{1},\ldots,u_{n}$ in (2.22) is a set of independent non-proper variables.
@@
Let us denote by $\mathcal{C}$ the set of the polynomial maps $F=H(u_{0},u_{1},\dots,u_{n}):\mathbb{K}^{2}\to\mathbb{K}^{2}$ with $u_{0},u_{1},\dots,u_{n}$ defined in (2.11); and by $\mathcal{C}^{n}_{r,s}$ we denote the corresponding subset of $\mathcal{C}$ for each fixed triple of natural numbers $(r,s,n)$. Theorems 2.18 and 2.20 say that the subsets $\mathcal{C}^{n}_{1,s}$ ($n\geq 1$, $s\geq 1$) and $\mathcal{C}^{1}_{r,s}$ ($r\geq 1$, $s\geq 1$) are the ones satisfying the Jacobian conjecture. Besides, from Proposition 2.23, we have:
@@
If a polynomial map belonging to a class $\mathcal{C}^{n}_{r,s}$ ($n\geq 2,r\geq 2,s\geq 1$) satisfies the Jacobian condition then it is a counterexample to the Jacobian conjecture.
@@ 3. Two-dimensional complex Jacobian conjecture of degree until 104
We consider now the complex case, i.e. consider polynomial maps $F=(F_{1},F_{2}):\mathbb{C}^{2}\to\mathbb{C}^{2}$. In this section, we will prove that the complex Jacobian conjecture is true for the degree until 104 (Theorem 3.6). For the case of degree 105, we reduce the problem into only four cases to consider (Proposition 3.7).
Let us recall the principal results concerning the complex Jacobian conjecture under the view point of Newton polygons studied by Abhyankar [2], Nagata [17, 18], Appelgate and Onishi [3], Magnus [13], Nakai and Baba [16] and Moh [14]. One can also find the results cited below in the book of van den Essen [8, pages 254-256].
@@
If ${\rm deg}\,(F_{1})$ divide ${\rm deg}\,(F_{2})$ or ${\rm deg}\,(F_{2})$ divide
${\rm deg}\,(F_{1})$, then the conjecture is true.
@@
Using so-called Newton-Puiseux chart, Henryk Żoła̧dek [24] give a stronger result than the one in Theorem 3.2:
@@
The Jacobian conjecture satisfies for maps with ${\rm gcd}({\rm deg}\,(F_{1}),{\rm deg}\,(F_{2}))\leq 16$ and for maps with ${\rm gcd}({\rm deg}\,(F_{1}),{\rm deg}\,(F_{2}))$ equal to 2 times a prime.
@@
The Jacobian conjecture satisfies for ${\rm deg}\,(F)\leq 104$.
@@
2) The case ${\rm deg}\,(F)=102$: Assume that ${\rm deg}\,(F_{1})<102$ and ${\rm deg}\,(F_{2})=102$. We see that $102=6.17$ and as $17$ is prime then if ${\rm deg}\,(F_{1})$ is not divisible by 17, the greatest common divisor of ${\rm deg}\,(F_{1})$ and ${\rm deg}\,(F_{2})$ is $d\leq 6$, and the conjecture is true by Theorem 3.2. Then we need to consider only the case ${\rm deg}\,(F_{1})$ is divisible by 17, i.e. the cases ${\rm deg}(F_{1})$ is 17, 34, 51, 68 and 85. If the degree of $F_{1}$ is 17, 34, 51 and 85, then by Theorem 3.3, the conjecture is true since ${\rm deg}\,(F_{1})$ is a product of at most two prime numbers. Now if ${\rm deg}\,(F_{1})=68$, then the greatest common divisor of ${\rm deg}\,(F_{1})$ and ${\rm deg}\,(F_{2})$ is $34$, which is two times a prime, therefore the conjecture is true by Theorem 3.5.
@@
If the conjecture satisfies for the cases where $({\rm deg}\,(F_{1}),{\rm deg}\,(F_{2}))$ is (42, 105), (63, 105), (70, 105) and (84, 105) then the conjecture satisfies for polynomial maps of degree 105.
@@
Assume that ${\rm deg}\,(F_{1})<105$ and ${\rm deg}\,(F_{2})=105$. We see that $105=3.5.7$ and as $7$ is prime then if ${\rm deg}\,(F_{1})$ is not divisible by 7, the greatest common divisor of ${\rm deg}\,(F_{1})$ and ${\rm deg}\,(F_{2})$ is $d\leq 15$, and the conjecture is true by Theorem 3.5. Then we need to consider only the case ${\rm deg}\,(F_{1})$ is divisible by 7. We remark that by Theorem 3.1 and Theorem 3.3, the cases we need to consider for ${\rm deg}\,(F_{1})$ are only: 7.4, 7.6, 7.8, 7.9, 7.10, 7.12, 7.14. The cases 7.4, 7.8 and 7.14 satisfy the conjecture by Theorem 3.2 since the greatest common divisor of ${\rm deg}\,(F_{1})$ and ${\rm deg}\,(F_{2})$ in these cases is a prime. So the remain cases are those for which the degree of $F_{1}$ is 42, 63, 70 and 84. ∎
@@
Theorems 2.18, Corollary 2.19 and Theorem 2.20 show that the complex conjecture is true for the constructed classes $\mathcal{C}^{n}_{1,s}$ ($n\geq 1$, $s\geq 1$) and $\mathcal{C}^{1}_{r,s}$ ($r\geq 2$, $s\geq 1$) (see also Remark 2.24). Then for studying the conjecture in our constructed class $\mathcal{C}$, we need to study only the maps belonging to the classes $\mathcal{C}^{n}_{r,s}$ ($n\geq 2,r\geq 2,s\geq 1$) (Corollary 2.25). Using non-proper variables of these classes, we can reduce the degree of the map. For example, if $n=2$, $r=2$ and $s=1$, then the non-proper variables in (2.11) are:

### v5

@@ Abstract.
We construct a non-proper set of two variables polynomial maps and study the nowhere vanishing Jacobian condition of the Jacobian conjecture for this set. We obtain some classes of polynomial maps satisfying the 2-dimensional Jacobian conjecture for both real and complex cases. In addition, by Newton polygon technique, we prove that the complex conjecture is true until degree 104, improving Moh boundary (degree 100) since 1983.
@@
Jacobian conjecture, polynomial maps, non-proper maps, Newton polygon
@@ 1. Introduction
([26, Lemma 2.2]). Then ${\rm det}(J(p,q)(x,y))>0$ for every $(x,y)\in\mathbb{R}^{2}$ since if $t=0$ then $f=y=\frac{1}{x}\neq 0$. However, $(p,q)$ is not automorphism. Our first observation is: by changing variables $t:=t(x,y)$, $h:=h(x,y)$ and $f:=f(x,y)$, we can write the Pinchuk maps $(p,q)$ as polynomial maps of new variables $t,h$ and $f$. The sequence $\{z_{k}\}=\{(k,1/k)\}\subset\mathbb{R}^{2}$ tends to infinity but $t(z_{k})=h(z_{k})=0$ and $f(z_{k})=1/k$. Then the Pinchuk maps $(p,q)$, being polynomial maps of new variables $t,h$ and $f$, satisfy that $(p(z_{k}),q(z_{k}))$ tends to a finite point and therefore they are non-proper and hence, is not automorphism. We call the set of such new variables $\{t,h,f\}$ a set of non-proper variables of the map $F=(p,q)$. By studying the Jacobian condition for polynomial maps of these new variables and classifying them, we obtain some classes that satisfy the Jacobian conjecture for both real and complex cases. The idea of “non-proper variables” first appeared in [23], where we called them “permanent variables”. We have used this idea to classify and stratify the non-proper set $S_{F}$ in order to calculate the intersection homology of singular varieties constructed in [22] (see also in [21] and [24]). This is another our approach to the Jacobian conjecture, by intersection homology. In this article, we also use Newton polygon technique in [2, 18, 19, 4, 15, 20, 27] to prove that the 2-dimensional complex Jacobian conjecture is true until degree 104, improving Moh boundary [16] (1983) from 100 till 104.
We would like to remark that in this article we try to provide another point of view to investigate the Jacobian conjecture using algebraical geometrical and topological approach, but there have been many approaches to the conjecture. One of the newest approaches is via quantization [3, 5], relating the Jacobian conjecture with the Dixmier conjecture.
@@
The first part, presented in Section 2, concerns the construction of non-proper variables for both real and complex cases and a classification of non-proper maps under the Jacobian condition. We obtain some explicit classes satisfying the 2-dimensional Jacobian conjecture for both real and complex cases (Corollaries 2.19, 2.21 and 2.22). These results are classified in Remark 2.24. We also provide a possible counter-example for the conjecture (if there exists) in Corollary 2.25.
The second part, presented in Section 3, concerns a particular result to the complex conjecture for degree until 104 (Theorem 3.6). We prove that the complex conjecture is true for degree until 104 using known results studied by Abhyankar [2], Nagata [18, 19], Appelgate and Onishi [4], Magnus [15], Nakai and Baba [20] and Żoła̧dek [27]. It is known that the complex conjecture is true for degree until 100, proved by Moh [16] in 1983 with a highly non-trivial proof using computer assistant improvement. There are also some works increasing the degree 100 of Moh, for instance [12], but we would like to emphasize that our proof is not a very difficult one, that may be surprising for this mysterious conjecture.
Let us remark that there may be a connection between the two parts of the paper to study the complex conjecture by this direction for higher degree cases (Remark 3.8).
@@ 2.1. Construction and examples
We consider $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$ and $F=(F_{1},F_{2}):\mathbb{K}^{2}\to\mathbb{K}^{2}$ a polynomial map of two variables $x$ and $y$. With $z\in\mathbb{K}^{2}$, we denote by $\|z\|$ its Euclidean norm. Let us recall firstly that the polynomial map $F$ is non-proper if there is a sequence $\{z_{k}=(x_{k},y_{k})\}_{k\in\mathbb{N}}\subset\mathbb{K}^{2}$ such that $\lim_{k\to\infty}\|z_{k}\|=+\infty$ and $\lim_{k\to\infty}F(z_{k})=(c_{1},c_{2})$, for some finite $c_{1},c_{2}\in\mathbb{K}$. The following remark is important for the definition of non-proper variables sets that we will introduce in Definition 2.2.
@@
In fact, if $F=(F_{1},F_{2})$ is non-proper, we can take $u_{0}:=F_{1}$ and $u_{1}:=F_{2}$ and the inverse is obvious. We define:
@@
then $\{u_{0},u_{1},u_{2},u_{3}\}$ is a set of non-proper variables of $F$ since any sequence $\{(1/k,-k+c)\}$, with $c\in\mathbb{C}$, satisfies the first condition in Definition 2.2. The second condition is verified by taking
@@
Now we show that $\{u_{0},u_{1},u_{2},u_{3}\}$ is independent. In fact, for any $z=(x,y)\in\mathbb{C}^{2}$, we have
@@
In the following example we will show that the functions $t,h,f$ defined in [26] and [11] form a set of non-proper variables for the Pinchuk maps and it is independent. Let us remark that we will consider both of real and complex cases, i.e. we consider the polynomials $t,h,f:\mathbb{K}^{2}\to\mathbb{K}$, with $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. In the complex case, it shows that the inverse of Corollary 2.8 is not true.
@@
    In fact, from [26, Lemma 2.1], it is enough to show that there is no $(x,y)\in\mathbb{K}^{2}$ such that
@@
    In fact, as in [26, page 3], set $p=f+h$, then by the Lemma 2.2 in [26], there exists a polynomial $u(f,h)$ such that
@@
As the remark at the end of the paper [26], the Pinchuk map constructed in the proof of [26, Lemma 2.2] has degree 40, where ${\rm deg}(p)=10$ and ${\rm deg}(q)=40$, but one can reduce ${\rm deg}(q)$ to 35 by a suitable choice of the auxiliary polynomial $u(f,h)$. In fact, the polynomial $q$ varies for different Pinchuk map, but always has the form $q=-t^{2}-6th(h+1)-u(f,h)$. In [10] (page 241), van den Essen choose
@@
In [6, 11], basing on Pinchuk’s examples [26], the authors provides new counterexamples to the real conjecture but with the possibility to reduce the degree of $q$ to 15: In [11, Corollary 2.6], they prove that if taking $p=f+h$ and
@@
(see the proof of [11, Proposition 2.3]). It follows:
@@ 2.2. A set of non-proper variables
has Newton polygon containing two points (1, 0) and (0, 1) (see Definition of Newton polygon in [2], [18, 19], also in [10, page 93]). In the following we will prove first that for any triple of non-zero natural numbers $(n,r,s)$ and for any polynomial mapping $F:\mathbb{K}^{2}\to\mathbb{K}^{2}$ such that
@@
where $H=(H_{1},H_{2}):\mathbb{K}^{n+1}\to\mathbb{K}^{2}$, then $\{u_{0},u_{1},\dots,u_{n}\}$ is a set of non-proper variables of $F$ (Proposition 2.12 for complex case and Proposition 2.13 for real case). Then we investigate the independence (Definition 2.4) for this set: we consider all the possibilities of triples $(r,s,n)$, we divide them into three cases: $r=1$ (section 2.2.2); $(n=1,\,r\geq 2)$ (section 2.2.3); and $(n\geq 2,r\geq 2)$ (section 2.2.4). We obtain some classes satisfying the Jacobian conjecture for polynomial mappings $F:\mathbb{K}^{2}\to\mathbb{K}^{2}$, both real and complex cases $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$ (Corollary 2.21 and Corollary 2.22).
@@
In the real case, if $s$ is even then the sequence $\{(x_{k},y_{k})\}$ chosen in the above proof can be used only if $b>0$. Therefore we have:
@@
If $s$ is odd, then the set of variables $u_{0},u_{1},\ldots,u_{n}:\mathbb{R}^{2}\to\mathbb{R}$ in (2.11) is a set of non-proper variables of any polynomial map $F(x,y)=H(u_{0},u_{,}\ldots,u_{n}):\mathbb{R}^{2}\to\mathbb{R}^{2}$ where $H=(H_{1},H_{2}):\mathbb{R}^{n+1}\to\mathbb{R}^{2}$ is a polynomial map of variables $u_{0},u_{1},\dots,u_{n}$. If $s$ is even, then the result holds when $b>0$.
@@
Considering the variables $u_{0},u_{1},\dots,u_{n}$ in (2.11), from Remark 2.3, we have the following corollaries:
@@
If $F:\mathbb{R}^{2}\to\mathbb{R}^{2}$ can be written as a polynomial map of variables $u_{0},u_{1},\dots,u_{n}$, i.e. $F(x,y)=H(u_{0},u_{1},\ldots,u_{n})$, where $H=(H_{1},H_{2}):\mathbb{R}^{n+1}\to\mathbb{R}^{2}$ is a polynomial map of variables $u_{0},u_{1},\dots,u_{n}$ in (2.11) then $F$ is non-proper if $s$ is odd. If $s$ is even, then the result holds when $b>0$.
@@
Then any point $z=(x_{0},\sqrt[s]{1/b})$, where $x_{0}\in\mathbb{K}$, is a solution of the system $\dfrac{\partial u_{i}}{\partial x}(z)=0$, $i=0,1,\ldots,n$ (notice that we consider $b\neq 0$ if $\mathbb{K}=\mathbb{C}$ and $b>0$ if $\mathbb{K}=\mathbb{R}$). ∎
@@
    In the real case $u_{0},u_{1}:\mathbb{R}^{2}\to\mathbb{R}$, the set of non-proper variables $\{u_{0},u_{1}\}$ is dependent with any $r$ odd, $s$ even and $b>0$.
@@
We have ${\partial u_{1}}/{\partial x}=1-rbx^{r-1}y^{s}$. With the same argument as in the proof of Theorem 2.18, the equation ${\partial u_{1}}/{\partial x}=0$ always has a solution in the complex case. In the real case, this equation has solution when $r$ is odd, $s$ is even and $b>0$. ∎
From Proposition 2.6, Corollary 2.19, Theorem 2.20 and Corollary 2.19, we have two explicit classes in which the two-dimensional Jacobian conjecture is true, for complex and real cases, respectively:
@@
$u_{0}=y,\quad u_{1}=x-bx^{r}y^{s}\quad(b\neq 0),$
@@
$u_{0}=y,\quad u_{1}=x-bx^{r}y^{s}\quad(r\,{\rm odd},\,s\,{\rm even},\,b>0),$
@@
For a fixed natural number $n\geq 2$, the set of variables $u_{0},u_{1},\ldots,u_{n}$ in (2.22) is a set of independent non-proper variables.
@@
Let us denote by $\mathcal{C}$ the set of the non-proper polynomial maps $F=H(u_{0},u_{1},\dots,u_{n}):\mathbb{K}^{2}\to\mathbb{K}^{2}$ with $u_{0},u_{1},\dots,u_{n}$ defined in (2.11) and by $\mathcal{C}^{n}_{r,s}$ the corresponding subset of $\mathcal{C}$ for each fixed triple of natural numbers $(r,s,n)$. The theorems 2.18 and 2.20 say that the classes $\mathcal{C}^{n}_{1,s}$ ($n\geq 1$, $s\geq 1$) and $\mathcal{C}^{1}_{r,s}$ ($r\geq 1$, $s\geq 1$) satisfy the complex Jacobian conjecture. This fact holds for the real conjecture if $r$ odd, $s$ even and $b>0$.
Moreover, from Proposition 2.23 and Remark 2.8, we have:
@@
A polynomial map belonging to a class $\mathcal{C}^{n}_{r,s}$ ($n\geq 2,r\geq 2,s\geq 1$) satisfying the Jacobian condition is a counterexample to the Jacobian conjecture.
@@ 3. Two-dimensional complex Jacobian conjecture of degree until 104
We consider now the complex case, i.e. considering polynomial maps $F=(F_{1},F_{2}):\mathbb{C}^{2}\to\mathbb{C}^{2}$. In this section, we will prove that the complex Jacobian conjecture is true for the degree until 104 (Theorem 3.6). For the case of degree 105, we reduce the problem into only four cases (Proposition 3.7).
Let us recall the principal results concerning the complex Jacobian conjecture under the view of point of Newton polygons studied by Abhyankar [2], Nagata [18, 19], Appelgate and Onishi [4], Magnus [15], Nakai and Baba [20] and Moh [16]. One can also find the results cited below in the book of van den Essen [10, pages 254-256].
@@
If ${\rm deg}\,(F_{1})$ divide ${\rm deg}\,(F_{2})$ or ${\rm deg}\,(F_{2})$ divide ${\rm deg}\,(F_{1})$, then the conjecture is true.
@@
Using so-called Newton-Puiseux chart, Henryk Żoła̧dek [27] give a stronger result than Theorem 3.2:
@@
The Jacobian conjecture satisfies for maps with
${\rm gcd}({\rm deg}\,(F_{1}),{\rm deg}\,(F_{2}))\leq 16$ and for maps with ${\rm gcd}({\rm deg}\,(F_{1}),{\rm deg}\,(F_{2}))$ equal to 2 times a prime.
@@
The 2-dimensional complex Jacobian conjecture satisfies for ${\rm deg}\,(F)\leq 104$.
@@
2) The case ${\rm deg}\,(F)=102$: Assume that ${\rm deg}\,(F_{1})<102$ and ${\rm deg}\,(F_{2})=102$. We see that $102=6.17$ and as $17$ is prime then if ${\rm deg}\,(F_{1})$ is not divisible by 17, the greatest common divisor of ${\rm deg}\,(F_{1})$ and ${\rm deg}\,(F_{2})$ is $d\leq 6$, and the conjecture is true by Theorem 3.2. Then we need to consider only the case where ${\rm deg}\,(F_{1})$ is divisible by 17, i.e. the cases ${\rm deg}(F_{1})$ is 17, 34, 51, 68 and 85. If ${\rm deg}(F_{1})$ is 17, 34, 51 and 85, then by Theorem 3.3, the conjecture is true since ${\rm deg}\,(F_{1})$ is a product of at most two prime numbers. Now if ${\rm deg}\,(F_{1})=68$, then the greatest common divisor of ${\rm deg}\,(F_{1})$ and ${\rm deg}\,(F_{2})$ is $34$, which is two times a prime (17 times 2), therefore the conjecture is true by Theorem 3.5.
@@
If the two-dimensional complex Jacobian conjecture satisfies for the four cases
$({\rm deg}\,(F_{1}),{\rm deg}\,(F_{2}))$ is (42, 105), (63, 105), (70, 105) and (84, 105) then the conjecture satisfies for polynomial maps of degree 105.
@@
Assume that ${\rm deg}\,(F_{1})<105$ and ${\rm deg}\,(F_{2})=105$. We see that $105=3.5.7$ and as $7$ is prime then if ${\rm deg}\,(F_{1})$ is not divisible by 7, the greatest common divisor of ${\rm deg}\,(F_{1})$ and ${\rm deg}\,(F_{2})$ is $d\leq 15$, and the conjecture is true by Theorem 3.5. Then we need to consider only the case ${\rm deg}\,(F_{1})$ is divisible by 7. We remark that by Theorem 3.1 and Theorem 3.3, the cases we need to consider for ${\rm deg}\,(F_{1})$ are only: 7.4, 7.6, 7.8, 7.9, 7.10, 7.12, 7.14. The cases ${\rm deg}\,(F_{1})=7.4,7.8$ and 7.14 satisfy the conjecture by Theorem 3.2 since the greatest common divisor of ${\rm deg}\,(F_{1})$ and ${\rm deg}\,(F_{2})$ in these cases is a prime. So the remain cases are those where ${\rm deg}\,(F_{1})$ is 42, 63, 70 and 84. ∎
To end the paper, we would like to mention the two following remarks:
@@
Theorem 2.18, Corollary 2.19 and Theorem 2.20 show that the complex conjecture is true for the constructed classes $\mathcal{C}^{n}_{1,s}$ ($n\geq 1$, $s\geq 1$) and $\mathcal{C}^{1}_{r,s}$ ($r\geq 1$, $s\geq 1$) (see also Remark 2.24). Then for the constructed class $\mathcal{C}=\cup\mathcal{C}^{n}_{r,s}$, we need to study the conjecture only for the maps belonging to the classes $\mathcal{C}^{n}_{r,s}$ ($n\geq 2,r\geq 2,s\geq 1$) (Corollary 2.25). Using non-proper variables of these classes, one can reduce the degree of the map. For example, if $n=2$, $r=2$ and $s=1$, then the non-proper variables in (2.11) are: